OpenStudy (anonymous):

RUSH! Determine whether the following sequence is arithmetic. If it is, find the common difference. If the sequence is not arithmetic, enter NS as the common difference. 10, 7, 4, 1, -2. -5 ...

OpenStudy (anonymous):

first term - second term = 10 -7 = 3 second term - third term = 7 -4 = 3 So it is AP: Start = 10, difff = -3 (MINUS 3)

OpenStudy (anonymous):

Find out the differences between every two consecutive terms. If they are the same, then its A.P. and that difference is the common difference.

OpenStudy (anonymous):

can i have the answer? im really sorry, im in rushhhh :((((9 :''( SOORRy

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

correct. THANK UUUU :D

OpenStudy (anonymous):

:-)